# Inference for Regression with

### v2

@@ 1 Introduction
As the amount of digitally recorded unstructured data continues to grow rapidly, empirical work in economics is increasingly incorporating it. The leading example of such data is text (Gentzkow et al., 2019a , Ash and Hansen, 2023), but others include surveys, images, and audio recordings. One of the primary applications of such data is to recover some latent variable of economic interest with an information retrieval (IR) model. Examples abound: Baker et al., (2016) measures economic policy uncertainty with newspaper text; Hoberg and Phillips, (2016) infers firms’ latent industries with corporate filings; Hansen et al., (2018) constructs measures of policy deliberation from Federal Open Market Committee (FOMC) transcripts; Magnolfi et al., (2022) uses survey data to measure product differentiation; Compiani et al., (2023) measures substitutability between products using Amazon text and image data; Gorodnichenko et al., (2023) measures tone-of-voice from audio recordings of FOMC press conferences; Gabaix et al., (2023) imputes firm characteristics from investor holdings data; Einav et al., (2022) infers patients’ health status from survey answers; Vafa et al., (2023) constructs a measure of labor market experience based on CVs. These derived measures are rarely an end in themselves. Rather, the motivation for constructing them is to study how the concept they proxy interacts with the economic environment. As such, they are typically plugged into downstream econometric models whose parameters are the main object of study. Importantly, the IR and econometric models are almost always treated as wholly separate, with the output of the former treated as “data” in the latter. We call this the two-step strategy.
@@
While clearly a pragmatic initial approach, the two-step strategy has largely unknown statistical properties. On one hand, ignoring the upstream IR model in the downstream econometric model suggests a generated regressor problem (Pagan, 1984). On the other, results in the time-series literature suggest plugging-in estimated latent variables need not lead to inference problems (Stock and Watson, 2002, Bernanke et al., 2005, Bai and Ng, 2006). More generally, characterizing the statistical guarantees—or lack thereof—of the two-step strategy is an important step in establishing a more mature understanding of reliable inference methods for unstructured data, an area that is still in its infancy.
@@
Our first contribution is to provide theoretical arguments for why the two-step strategy leads to biased inference on regression parameters in empirically plausible settings. We consider a set of $n$ observations of quantitative and unstructured data. Each unstructured observation is composed of a high-dimensional vector of feature counts, [^1] For example, one of the simplest representations of a textual corpus is the bag-of-words model in which each document is represented as a vector of integer counts over the unique vocabulary terms in the corpus. Even relatively small corpora contain thousands of unique dimensions. Moreover, the dimensionality grows even further as one consider richer linguistic units than individual words. More generally, many unstructured datasets can be represented as high-dimensional, categorical data. where $C_{i}$ is the amount of unstructured data for observation $i$. The relative magnitudes of $n$ versus moments of $C_{i}$ play a key role in our analysis. We next specify a statistical model with three parts: a distribution over the feature-count vectors; a low-dimensional, latent variable representation for each such distribution; and a regression of an observed outcome variable onto the latent variables. The two-step strategy (i) estimates the latent variables from the observed feature counts, then (ii) regresses the outcome variable onto these estimates. This procedure mimics the common approach in the empirical literature described above. Our primary theoretical question is: under what conditions do the estimated coefficients and standard errors from (ii) allow for valid inference?
@@
The basic problem is measurement error: the regressors in step (ii) contain estimated rather than true latent variables. As is well known, measurement error leads to biased point estimates and distorted standard errors, both of which are present as the number of observations $n$ grows with a fixed amount of unstructured data per observation. To capture a more empirically realistic situation, we allow $n$ and the distribution of $C_{i}$ to grow together so that both sampling error and measurement error are relevant for inference. [^2] This framework is motivated by the observation that both sample size and the amount of unstructured data per observation are typically large in applications. Conventional fixed-DGP asymptotics provide poor approximations to this case. Our use of sequences of DGPs to get better asymptotic approximations to finite-sample behavior is similar in spirit to the weak instrument literature (Staiger and Stock, 1997), earlier work on measurement error (Chesher, 1991), and unit root testing (Phillips, 1987). Our main finding is that, whenever $\sqrt{n}\times\mathbb{E}\left[\frac{1}{C_{i}}\right]$ tends to a constant $\kappa>0$, there is a bias present in the asymptotic distribution of the regression coefficients which is increasing in $\kappa$. Larger values of $\kappa$ give relatively greater importance to measurement error, and hence a larger bias. However, the asymptotic variance is the same as that from regression onto the true latent variables and the usual OLS standard errors are consistent. Hence, treating the estimated latent variables as observed data in the regression does not distort the width of confidence intervals, but centers them away from the truth. This contrasts with the generated regressor literature, which emphasizes the variance distortion arising from treating plug-in estimates as data. [^3] In the classic generated regressor problem (Pagan, 1984) there is a common finite-dimensional parameter estimated in the first stage whereas here all $n$ latent covariates are estimated. Only when $\kappa=0$, so that sampling error dominates measurement error, does the two-step strategy allow for valid inference.
@@
Of course, $\kappa$ describes limiting behavior and cannot be used to directly compute the magnitude of the bias in a given finite dataset. But our theoretical arguments provide insight into when this bias is potentially problematic. Take, for example, job postings data as recorded by Lightcast (formerly Burning Glass), which has been used in dozens of papers. In 2022, there were 45 million individual job postings in the United States, with an average inverse posting length of 0.003. The empirical analogue of $\kappa$ is $\sqrt{45,000,000}\times 0.003\approx 20$, suggesting that measurement error may be large relative to sampling error. Another insight of the theory is that the magnitude of the bias arises not from the average amount of unstructured data per observation but rather the average inverse amount. So, if a dataset has a long tail of observations with little data, a bias can arise even if there is a substantial amount of data per observation on average. Again taking the Lightcast data, the average document size is $575$. If one used $1/575$ in place of $\mathbb{E}\left[C^{-1}_{i}\right]$ to compute the analogue of $\kappa$, $20$ would fall to below $12$, highlighting the role of this tail behavior in driving measurement error bias. Nor is the Lightcast data a special case. There were $315,000$ patents filed to the US Patent and Trademark Office in 2023 and their summary texts have an average inverse document length of 0.002, so the analogue of $\kappa$ is above $1$. The calculation based instead on the inverse of the average document length is $0.4$. Of course, there are other cases where one may have few total observations but with each individual observation having a large amount of unstructured data. [^4] This case is similar to that typically considered in the literature on factor-augmented regression, which extracts common factors from time series and plugs them into downstream regressions. There are typically dozens or hundreds of time series $N$ per time unit, but limited observations $T$ per series. In our setting, $T$ (respectively $1/N$) is analogous to $n$ (respectively $\mathbb{E}[C^{-1}_{i}]$). Bai and Ng, (2006) show that factor augmentation leads to valid inference when $\sqrt{T}/N\rightarrow 0$, analogous to $\kappa=0$. Because the exact magnitude of the problem is hard to assess in any given setting, it is important to develop robust inference methods that guard against measurement error whenever it may be present, but still allow reliable inference when it is not.
@@
While implementing the one-step strategy is straightforward theoretically, it presents a major computational challenge due to the large number of observation-specific latent variables that must be integrated out. To address this, we use Hamiltonian Monte Carlo (HMC; MacKay, 2003, Neal, 2012), a Markov Chain Monte Carlo algorithm that uses information on the gradient of a distribution to sample from it. Implementation is greatly simplified with the use of modern probabilistic programming languages: one simply specifies the likelihood in code, which is then “automatically” compiled to perform sampling. This paradigm is useful for applied researchers because it allows one to focus on model development without the need to re-write the estimation and inference algorithms each time the model is changed. [^5] Previous papers that have performed inference using the joint likelihood approach with unstructured data include Gentzkow et al., 2019b , Ruiz et al., (2020), and Munro and Ng, (2022). These typically require extensive code to estimate, which makes adapting the model difficult for non-specialists. The most common probabilistic programming language in applied econometrics is Stan (Carpenter et al., 2017), which has been used by, for example, Meager, (2019) and Bandiera et al., (2021). Such applications have been limited to simple Bayesian meta-analyses with a few dozen parameters. Instead, we use the NumPyro package (Bingham et al., 2018, Phan et al., 2019), which efficiently computes the gradients that underlie HMC by using massive parallelization on dedicated hardware.
Third, we compare the performance of the two-step and one-step strategies in an applied setting. To this end, we introduce the Supervised Topic Model with Covariates (STMC) which combines elements of existing models (Blei et al., 2003, Roberts et al., 2014, Ahrens et al., 2021) but is, to the best of our knowledge, a new statistical model of unstructured data. The model reduces the dimensionality of feature-count vectors by projecting them onto a set of latent factors (or topics), as in Probabilistic Latent Semantic Analysis (Hofmann, 1999) and Latent Dirichlet Allocation (Blei et al., 2003). The dependence of outcome variables on latent factor loadings and observed covariates is captured by a “downstream” regression model. Additionally, the factor loadings can depend on a potentially different set of covariates via a second “upstream” regression model. All components are woven together by a joint likelihood. Specifying the model in code takes fewer than 25 lines, illustrating how one can perform automatic inference in a new setting that would previously have required a bespoke and complex codebase.
@@
In simulated data, we show that the two-step strategy produces estimates that exhibit a bias which is increasing in $\kappa$. Moreover, two-step confidence interval widths are similar to those obtained using the true latent variables as covariates. Both of these findings reinforce the main predictions of our theory. By contrast, the one-step strategy produces estimates that appear unbiased and the corresponding confidence intervals have the same widths as those using the true latent variables. Thus, one-step confidence intervals have both the correct width and the correct centering.
@@
Next, we revisit the empirical application from Bandiera et al., (2020) which uses the two-step strategy to first estimate latent CEO behaviors from a CEO time-use survey, then explains firm performance using the estimated behaviors. The one-step strategy substantively changes estimates compared to the two-step strategy. For instance, the estimated effect of having an MBA degree on behavior more than doubles depending on whether the one-step or two-step strategy is used. To further test our theory, we next reduce the amount of unstructured data per observation and again deploy both inference strategies. This increases measurement error in latent behavior, and hence should increase the bias of the two-step strategy. Since the one-step strategy is always unbiased (asymptotically), one should observe larger differences in estimates, which is what we find. The estimated impact of behavior on firm performance, equivalent for both methods in the original data, is now over twice as large under the one-step strategy. Moreover, under the one-step strategy, the estimated effect sizes on behavior of having an MBA degree and of managing a large firm now triple.
Our overall message is that a popular way of using unstructured data in empirical work may suffer from measurement error which biases inference. We are unaware of other papers that explicitly model the source of this error, and how it relates to sampling error. Ultimately it is this trade-off (as manifested in $\kappa$) that is most important for inference. [^6] Fong and Tyler, (2021), Allon et al., (2023), and Zhang et al., (2023) all assume the existence of measurement error in a supervised learning algorithm used to generate regressors, but do not tie it to a specific model. Their proposed solution relies on a correctly labeled subset of data which can be used to build IV/GMM estimators. The one-step strategy requires no such labeled dataset to produce unbiased estimates. It can also be extended easily to handle non-linear models with measurement error. On a more positive note, though, a solution exists that is relatively easy to implement and computationally feasible. Since in practice researchers cannot characterize the severity of the measurement error in a given dataset, and there is little downside to applying the one-step strategy, we see it as a robust starting point for empirical analysis. We do note, however, that implementing the one-step strategy requires formulating a likelihood function. Latest-generation machine-learning- and AI-based approaches to information retrieval increasingly use neural networks with no obvious statistical structure that yields a likelihood function. A first comment is that, while implementing the one-step strategy may not be possible in these settings, the measurement error problem does not thereby disappear. Instead, it simply becomes harder to characterize statistically. Second, such approaches are often given statistical foundations following their adoption and, as this process plays out over the coming years, the scope for the one-step strategy will expand accordingly. [^7] One illustrative example is the popular word2vec model for producing word embeddings. The original model (Mikolov et al., 2013b , Mikolov et al., 2013a ) had no statistical interpretation but yielded word representations that nevertheless captured semantic relationships well. Word2vec has subsequently been adopted by economists as part of the two-step strategy, for example to measure occupation-level exposure to technological change (Kogan et al., 2019) and emotionality in political speech (Gennaro and Ash, 2022). In parallel, a literature has developed likelihood-based interpretations of embeddings (Arora et al., 2016, Dieng et al., 2020, Ruiz et al., 2020) which could in principle be adapted for use in the one-step strategy. More generally, our belief is that inference problems arising from the analysis of unstructured data should be better recognized and taken more seriously in order to fully harness its potential value.
@@ 2.1 Stylized Model
Policy uncertainty itself is a nebulous concept that is difficult to precisely define let alone observe. The key innovation of Baker et al., (2016) is to construct EPU indices based on monthly counts of articles in 10 newspapers containing certain terms, then convert to index form. Their EPU index is then introduced as a covariate in regressions and VARs. But it’s arguably the case that their measure, while a strong signal of policy uncertainty, is not numerically the same as policy uncertainty. For instance, one could change the set of newspapers surveyed and obtain a quantitatively different (but related) measure. We therefore adopt the specification

### v3

@@ 1 Introduction
The amount of unstructured data is growing rapidly and empirical work in economics is increasingly using it. The leading example of such data is text (Gentzkow et al., 2019a , Ash and Hansen, 2023), but others include surveys, images, and audio recordings. In economics, unstructured data is primarily used to measure variables of interest that aren’t observed in conventional quantitative data sources. Examples abound: Baker et al., (2016) measure economic policy uncertainty with newspaper text; Hoberg and Phillips, (2016) infer firms’ latent industries with corporate filings; Hansen et al., (2018) construct measures of policy deliberation from Federal Open Market Committee (FOMC) transcripts; Magnolfi et al., (2022) use survey data to measure product differentiation; Compiani et al., (2023) measure substitutability between products using Amazon text and image data; Gorodnichenko et al., (2023) measure tone-of-voice from audio recordings of FOMC press conferences; Gabaix et al., (2023) impute firm characteristics from investor holdings data; Einav et al., (2022) infer patients’ health status from surveys; Vafa et al., (2023) construct a measure of labor market experience based on CVs. These derived measures are rarely an end in themselves. Rather, the goal is to study how the latent variables they measure interact with the economic environment. As such, they are typically plugged-in to downstream econometric models whose parameters are the main object of study. In practice, the upstream information retrieval (IR) model used to extract measurements from unstructured data and the downstream econometric model are almost always taken as wholly separate: the output of the upstream model is treated as observed “data” in the downstream model. We call this the two-step strategy.
@@
While clearly a pragmatic initial approach, the two-step strategy has largely unknown statistical properties. On one hand, ignoring the upstream estimation step in downstream inference suggests a generated regressor problem (Pagan, 1984). On the other, results in the time-series literature suggest plugging-in estimated latent variables need not lead to inference problems (Stock and Watson, 2002, Bernanke et al., 2005, Bai and Ng, 2006). More generally, characterizing the statistical guarantees—or lack thereof—of the two-step strategy is an important step in establishing a more mature understanding of reliable inference methods for unstructured data, an area that is still in its infancy.
@@
Our first contribution is to provide theoretical arguments for why the dominant two-step strategy leads to biased inference on regression parameters in empirically plausible settings. We consider a set of $n$ observations of quantitative and unstructured data. As many unstructured datasets can be represented as high-dimensional categorical data, we treat each unstructured observation as a high-dimensional vector of feature counts. [^1] For example, one of the simplest representations of text is the bag-of-words model. Each document is represented as a vector of integer counts over the unique vocabulary terms in the corpus. Even relatively small corpora contain thousands of unique dimensions. We allow the amount of unstructured data to vary across observations, letting $C_{i}$ denote the amount of unstructured data for observation $i$. The relative magnitudes of $n$ versus moments of $C_{i}$ play a key role in our analysis. We next specify a statistical model with three parts: a distribution over the feature-count vectors; a low-dimensional, latent variable representation for each such distribution; and a regression of an observed outcome variable onto the latent variables. [^2] This model is the basis for a large empirical literature. Examples include Nimczik, (2017), Hansen et al., (2018), Mueller and Rauh, (2018), Larsen and Thorsrud, (2019), Bandiera et al., (2020), Thorsrud, (2020), Bybee et al., (2020), Adams et al., (2021), Draca and Schwarz, (2021), Olivella et al., (2021), Ash et al., (2022), and Munro and Ng, (2022). The two-step strategy (i) estimates the latent variables from the observed feature counts, then (ii) regresses the outcome variable on these estimates. Our primary theoretical question is: under what conditions do the estimated coefficients and standard errors from (ii) allow for valid inference?
@@
The basic problem is measurement error: the regressors in step (ii) contain estimated rather than true latent variables. As is well known, measurement error leads to inconsistent estimates and distorted standard errors, both of which are present as the number of observations $n$ grows with a fixed amount of unstructured data per observation. Because both the sample size and the amount of unstructured data per observation are typically large in applications, we allow $n$ and the distribution of $C_{i}$ to evolve together so that sampling error and measurement error are both relevant for inference. [^3] Our use of sequences of DGPs to better approximate the finite-sample behavior of estimators is similar in spirit to the weak instrument literature (Staiger and Stock, 1997), earlier work on measurement error (Chesher, 1991), unit root testing (Phillips, 1987), and large $n,T$ panels (Hahn and Kuersteiner, 2002). Our main finding is that, whenever $\sqrt{n}\times\mathbb{E}\left[\frac{1}{C_{i}}\right]$ tends to a constant $\kappa>0$, the estimator is consistent but there is a bias present in the asymptotic distribution of the regression coefficients. Larger values of $\kappa$ give relatively greater importance to measurement error, and hence a larger bias. However, the asymptotic variance is the same as that from regression onto the true latent variables and the usual OLS standard errors are consistent. Hence, treating the estimated latent variables as observed data in the regression does not distort the width of confidence intervals, but centers them away from the truth. This contrasts with the generated regressor literature, which emphasizes the variance distortion arising from plugging-in estimates as data. [^4] In the classic generated regressor problem (Pagan, 1984) there is a common finite-dimensional parameter estimated in the first stage whereas here all $n$ latent covariates are estimated. Only when $\kappa=0$, so that sampling error dominates measurement error asymptotically, does the two-step strategy allow for valid inference.
@@
Of course, $\kappa$ describes limiting behavior and cannot be used to directly compute the magnitude of the bias in a given finite dataset. But our theory provides insights into when this bias is potentially problematic. Take, for example, job postings data as recorded by Lightcast (formerly Burning Glass), which has been used in dozens of papers. In 2022, there were 45 million individual job postings in the United States, with an average inverse posting length of 0.003. The empirical analogue of $\kappa$ is $\sqrt{45,000,000}\times 0.003\approx 20$, suggesting that measurement error may be distorting inference. Another example is the popular Nielsen Homescan database, where the empirical analogue of $\kappa$ is approximately $4$. A second insight is that the magnitude of the bias arises from the average inverse amount of unstructured data per observation. So, if a dataset has a long tail of observations with little data, a bias can arise even if there is a substantial amount of undstructured data per observation on average. Again taking the Lightcast data, the average document size is $575$. If one used $1/575$ in place of $\mathbb{E}\left[C^{-1}_{i}\right]$ to compute the analogue of $\kappa$, $20$ would fall to below $12$, highlighting the role of tail behavior in driving measurement error bias. There are other cases, such as US patent data, where the empirical analogue of $\kappa$ takes a smaller value around 1 because each observation has a relatively large amount of unstructured data. Because the exact magnitude of the problem is hard to assess in any given setting, it is important to develop robust inference methods that correctly account for measurement error.
@@
While implementing the one-step strategy is straightforward theoretically, it presents a major computational challenge due to the large number of observation-specific latent variables that must be integrated out. To address this, we use Hamiltonian Monte Carlo (HMC; MacKay, 2003, Neal, 2012), a Markov Chain Monte Carlo algorithm that uses information on the gradient of a distribution to sample from it. Implementation is greatly simplified with the use of modern probabilistic programming languages: one simply specifies the likelihood in code, which is then “automatically” compiled to perform sampling. This paradigm is useful for applied researchers because it allows one to focus on model development without the need to re-write the estimation and inference algorithms each time the model is changed. To this end, we use the NumPyro package (Bingham et al., 2018, Phan et al., 2019), which requires only a few lines of code to implement and is well suited to handling models with a large number of latent variables. [^5] Previous papers that have performed inference using the joint likelihood approach with unstructured data include Gentzkow et al., 2019b , Ruiz et al., (2020), and Munro and Ng, (2022). These typically require extensive code to estimate, which makes adapting the model difficult for non-specialists.
Third, we compare the performance of the two-step and one-step strategies in an applied setting. To this end, we introduce the Supervised Topic Model with Covariates (STMC) which combines elements of existing models (Blei et al., 2003, Roberts et al., 2014, Ahrens et al., 2021) but is, to the best of our knowledge, a new statistical model of unstructured data. The model reduces the dimensionality of feature-count vectors by projecting them onto a set of latent factors (or topics), as in Probabilistic Latent Semantic Analysis (Hofmann, 1999) and Latent Dirichlet Allocation (Blei et al., 2003). The dependence of outcome variables on latent factor loadings and observed covariates is captured by a downstream regression model. Additionally, the factor loadings can depend on a potentially different set of covariates via the upstream model. All components are woven together by a joint likelihood. Specifying the model in code takes fewer than 25 lines, illustrating how one can perform automatic inference in a new setting that would previously have required a bespoke and complex codebase.
@@
In simulated data, we show that the two-step strategy produces estimates that exhibit a bias which is increasing in $\kappa$. Moreover, two-step confidence interval widths are similar to those obtained using the true latent variables as covariates. Both of these findings reinforce the main predictions of our theory. By contrast, the one-step strategy produces estimates that appear unbiased and confidence intervals that have both the correct width and the correct centering.
@@
Next, we revisit the empirical application from Bandiera et al., (2020) which uses the two-step strategy to first estimate latent CEO behaviors from a CEO time-use survey, then explains firm performance using the estimated behaviors. The one-step strategy substantively changes estimates compared to the two-step strategy. For instance, the estimated effect of having an MBA degree on behavior more than doubles when using the one-step instead of two-step strategy. To further test our theory, we next reduce the amount of unstructured data per observation and again deploy both inference strategies. This increases measurement error in latent behavior, and hence should increase the bias of the two-step strategy. Since the one-step strategy is always unbiased (asymptotically), one should observe larger differences in estimates, which is what we find. The estimated impact of behavior on firm performance, equivalent for both methods in the original data, is now half as large under the two-step strategy. Moreover, under the two-step strategy, the estimated effects of CEO characteristics on behavior reduce by a factor of three.
The settings we consider have an upstream unsupervised learning problem where unstructured data is used to recover latent variables of interest. A parallel set of recent papers considers the upstream generation of regressors via supervised learning (Fong and Tyler, 2021, Allon et al., 2023, Egami et al., 2023, Zhang et al., 2023). These papers rely on the strong assumption that the latent variable is observed with no error for a subset of observations. Their proposed solution uses the correctly labeled subset of data to build IV/GMM estimators. In our baseline setting, in common with much of the economics literature, all latent variables remain unobserved, but our framework could be extended easily to incorporate noisy labels.
Gentzkow et al., 2019b highlights measurement error in the estimation of group differences in observed speech. This error emerges when the number of observed draws per group (analogous to $C_{i}$ in our setting) is small relative to the number of possible vocabulary terms. To examine this situation, we extend our baseline model to consider regressions on text-derived similarity measures. Analogously to Gentzkow et al., 2019b and our baseline setting, there is no bias when average inverse document size grows sufficiently fast with the sample size $n$. But when it doesn’t, so that measurement error is also relevant asymptotically, a bias again emerges.
Our overall message is that the dominant approach for using unstructured data in empirical work in economics may suffer from measurement error which biases inference. We illustrate this formally with theoretical arguments in an empirically relevant albeit restrictive setting, but the take-away applies much more broadly. On a more positive note, a solution exists that is easy to implement and computationally feasible. We therefore see the one-step strategy as a robust and widely applicable starting point for empirical analysis. For instance, an emerging line of research uses text-derived sentiment indices as inputs into forecasting models with a vector autoregressive or dynamic factor structure. Straightforward extensions of our theoretical arguments can be used to show how error in the indices will bias coefficient estimates and limit the effectiveness of these forecasting methods. More constructively, the one-step strategy can be used to enhance the performance of these forecasting methods. Likewise, the industrial organization literature is increasingly using embedded representations of firms and products to characterize market behavior and demand with structural models. Our one-step strategy can be used to mitigate bias introduced by measurement error in the embeddings in these. Going forward, it is important to establish for which specific IR methods and econometric models does measurement error most severely affect inference. More generally, our belief is that inference problems arising from the analysis of unstructured data should be better recognized and taken more seriously in order to fully harness its potential value.
@@ 2.1 Stylized Model
Policy uncertainty itself is a nebulous concept that is difficult to precisely define let alone observe. The key innovation of Baker et al., (2016) is to construct EPU indices based on monthly counts of articles in ten newspapers containing certain terms, then convert to index form. Their EPU index is then introduced as a covariate in regressions and VARs. But it’s arguably the case that their measure, while a strong signal of policy uncertainty, is not numerically the same as policy uncertainty. For instance, one could change the set of newspapers surveyed and obtain a quantitatively different (but related) measure. We therefore adopt the specification